OpenStudy (anonymous):

will you please show me a graph -8, -11

OpenStudy (anonymous):

u mean graph that coordinate?

OpenStudy (anonymous):

(-8, -11)?

OpenStudy (anonymous):

It is just a point 8 units left of the x/y axis, 11 units down

OpenStudy (anonymous):

yepp :) agreed with rickjbr!

OpenStudy (anonymous):



OpenStudy (anonymous):

I have to show each system by graphing. its y = 3x =3 y = x-3 The answer is -8, and -11. But what does the graph look like to this I need to show the arrows

OpenStudy (anonymous):

You plug in values for x and solve for y. Then you draw a line through two or more points. You only need 2 points to draw the line, however.

OpenStudy (anonymous):

Therefore, Rick what does the graph look like

OpenStudy (anonymous):

Why do you say y=3x=3?

OpenStudy (anonymous):

Is there one equation or more?

OpenStudy (anonymous):

What is the equation you are trying to graph?

OpenStudy (anonymous):

No its y=3x+13 Y=x-3 im sorry

OpenStudy (anonymous):

You need to graph two lines?